OpenStudy (anonymous):

Which of the following constants can be added to x^2 - 10x to form a perfect square trinomial? 10 25 100

OpenStudy (anonymous):

a)10 b)25 c)100

OpenStudy (amistre64):

A perfect square trinomial is of the form: ax^2 +bx +c where c=(b/2)^2. x^2 -10x +c where c=(-10/2)^2 = (-5^2) = 25

OpenStudy (amistre64):

...also known as Completing the Square. Not a very cool sounding name, but informative nonetheless :)